# Multi-Timescale Gradient Sliding for Distributed Optimization

### v1

@@ Abstract
We propose two first-order methods for convex, non-smooth, distributed optimization problems, hereafter called Multi-Timescale Gradient Sliding (MT-GS) and its accelerated variant (AMT-GS). Our MT-GS and AMT-GS can take advantage of similarities between (local) objectives to reduce the communication rounds, are flexible so that different subsets (of agents) can communicate at different, user-picked rates, and are fully deterministic. These three desirable features are achieved through a block-decomposable primal-dual formulation, and a multi-timescale variant of the sliding method introduced in [lan_communication-efficient_2020, lan_gradient_2016], where different dual blocks are updated at potentially different rates.
To find an $\epsilon$-suboptimal solution, the complexities of our algorithms achieve optimal dependency on $\epsilon$: MT-GS needs $O(\overline{r}A/\epsilon)$ communication rounds and $O(\overline{r}/\epsilon^{2})$ subgradient steps for Lipchitz objectives, and AMT-GS needs $O(\overline{r}A/\sqrt{\epsilon\mu})$ communication rounds and $O(\overline{r}/(\epsilon\mu))$ subgradient steps if the objectives are also $\mu$-strongly convex. Here, $\overline{r}$ measures the “average rate of updates” for dual blocks, and $A$ measures similarities between (subgradients of) local functions. In addition, the linear dependency of communication rounds on $A$ is optimal [ArjevaniShamir2015Communication], thereby providing a positive answer to the open question whether such dependency is achievable for non-smooth objectives [ArjevaniShamir2015Communication].
@@ 1 Introduction
Distributed optimization is a branch of optimization, where multiple agents, each having access to only partial information about the (global) objective, work together to solve the global problem. As an example, in distributed empirical risk minimization for machine learning, the global objective function is the sum of local loss functions, each depending on the local dataset which is only available to one agent [AssranAytekinFeyzmahdavianJohanssonRabbat2020, Boyd2011, Zhu2016, ArjevaniShamir2015Communication, Jakub2016]. Examples of other applications include power system control [Molzahn2017, Nedic2018], multi-robot system control [Bullo2009, Nedic2018, Jaleel2020, Shorinwa2024], and signal processing [BertsekasTsitsiklis1989, Dan2002, Rabbat2004], to name a few.
In this work, we study convex, non-smooth, distributed optimization problems of the following form:
$\min_{x\in\mathcal{X}}\sum_{v\in V}f_{v}(x)$ ($\mathcal{P}$)
@@
where $V=[m]$ represents $m$ agents, $\mathcal{X}\subset\mathbb{R}^{d}$ is a nonempty, closed convex set, and $f_{v}:\mathcal{X}\to\mathbb{R}$ is a convex and possibly non-smooth objective function such that for some $M,\mu\geq 0$, we have for all $v\in V$,
$\frac{\mu}{2}\|x-y\|^{2}\leq f_{v}(x)-f_{v}(y)-\langle f^{\prime}_{v}(y),x-y\rangle\leq M\|x-y\|,\quad\forall x,y\in\mathcal{X},$ (1)
@@
where $f_{v}^{\prime}:\mathcal{X}\to\mathbb{R}^{d}$ is a subgradient oracle, i.e. $f^{\prime}_{v}(x)\in\partial f_{v}(x)$ for all $x\in\mathcal{X}$, and $f^{\prime}_{v}$ is only available to agent $v$. For instance, when $\|f^{\prime}_{v}\|_{*}\leq M_{f}$, $M=2M_{f}$ holds. [^1] This is because $f_{v}(y)\geq f_{v}(x)+\langle f^{\prime}_{v}(x),y-x\rangle$, and so using Cauchy–Schwarz inequality $f_{v}(x)-f_{v}(y)-\langle f^{\prime}_{v}(y),x-y\rangle\leq\langle f^{\prime}_{v}(x)-f^{\prime}_{v}(y),x-y\rangle\leq 2M_{f}\|x-y\|.$
We assume that agent $v$ maintains and updates $x_{v}$, a local version of the decision variable $x$. In this setting, due to the lack of “global views” from agents’ (local) perspectives, information aggregation – such as communication and averaging – is necessary for agents to reach consensus and approximately solve ($\mathcal{P}$).
We target settings where within and between different subsets of the local objectives, the scales of the function similarities and the costs of information aggregation could be (vastly) different. For instance, in distributed empirical risk minimization, local loss functions could inherit potential similarities in local datasets, which helps reduce the communication round needed [Zhu2016, Kovalev2022, pmlr-v32-shamir14, pmlr-v37-zhangb15, pmlr-v151-tian22b, ArjevaniShamir2015Communication]; the costs of communication could depend on factors such as the distance between agents, methods of communication, and amounts of data sent [Berahas2019, BertsekasTsitsiklis1989, Scaman2019, Tyurin2024, Santosh2020].
The heterogeneity in the function similarities and the communication costs makes it desirable to have more refined control over the numbers of communication rounds among different subsets of agents. In addition, sometimes stochastic algorithms are impractical or inefficient due to factors such as unpredictability, random memory access [Sun2017], and sampling overhead [Csiba2018]. For these reasons, we aim at designing deterministic algorithms which allow users to pick different numbers of communication rounds among different subsets of agents. Toward this end, we make contributions to both the problem formulation and the algorithm design.
We propose a saddle point reformulation of ($\mathcal{P}$) with block-decomposable duals, where each dual block imposes “consensus constraints” within a subset of the agents, and when a dual block is updated, information is aggregated within the corresponding agents. Taking advantage of the block-decomposable structure, we propose updating different dual blocks at different rates, which leads to the Multi-Timescale Gradient Sliding (MT-GS) algorithm and its accelerated variant (AMT-GS).
@@
To find an $\epsilon$-suboptimal solution, the complexities of our algorithms achieve optimal dependency on $\epsilon$: MT-GS needs $O(\overline{r}A/\epsilon)$ communication rounds and $O(\overline{r}/\epsilon^{2})$ subgradient steps for Lipchitz objectives, and AMT-GS needs $O(\overline{r}A/\sqrt{\epsilon\mu})$ communication rounds and $O(\overline{r}/(\epsilon\mu))$ subgradient steps if the objectives are also $\mu$-strongly convex. Here, $\overline{r}$ measures the “average rate of updates” for dual blocks, and $A$ measures similarities between (subgradients of) local functions. In addition, the linear dependency of communication rounds on $A$ is optimal [ArjevaniShamir2015Communication], thereby providing a positive answer to the open question whether such dependency is achievable for non-smooth objectives [ArjevaniShamir2015Communication].
@@
Numerical experiments for the support vector machine problem with regularized hinge losses confirm the effectiveness of our algorithms and demonstrate the above dependence on $\overline{r}$ and $A$.
@@ 1.1 Related works
Non-smooth distributed optimization. Since the seminal works [BertsekasTsitsiklis1989, TsitsiklisBertsekasAthans1986], numerous algorithms have been proposed for non-smooth distributed optimization under various settings, and we refer readers to surveys such as [AssranAytekinFeyzmahdavianJohanssonRabbat2020, NedicOlshevskyRabbat2018, konecny2016federatedoptimizationdistributedmachine]. For the function class of Lipschitz, non-smooth, convex objectives, most of these algorithms fall into the following two categories: subgradient based and dual based [lan_communication-efficient_2020]. Subgradient based algorithms such as the incremental gradient method [bertsekas_incremental_2011], decentralized subgradient method [NedicOzdaglar2009], and the dual averaging [Duchi2012DA] usually require $O(1/\epsilon^{2})$ rounds of communication, each followed by one gradient step. Within the function class, this achieves the optimal subgradient oracle complexity, but is suboptimal with respect to the communication rounds: as proven by [ArjevaniShamir2015Communication, Scaman2019], the communication rounds needed is $O(1/\epsilon)$.
@@
As a comparison, dual based algorithms, which dualize the consensus constraints, usually have better communication complexity: $O(1/\epsilon)$ rounds are needed for distributed ADMM [Aybat2018, wei2013o1kconvergenceasynchronousdistributed] and the decentralized communication sliding (DCS) [lan_communication-efficient_2020], as examples. However, each round of communication is followed by optimization of Lagrangians or proximal updates, performed locally by each agent. To make the overall algorithm first-order, [lan_communication-efficient_2020] proposes the Communication Sliding (CS) procedure, which approximates the proximal updates through $O(1/\epsilon)$ steps of (local) mirror descent, thereby achieving the $O(1/\epsilon^{2})$ subgradient oracle complexity. The CS procedure has roots in the gradient sliding technique [lan_gradient_2016], which can save gradient computation for the smooth component when the objective involves a smooth and a non-smooth component.
@@
For the class of strongly convex objectives, DCS can be accelerated, needing $O(1/\sqrt{\epsilon})$ rounds of communication and $O(1/\epsilon)$ gradient steps in total, both achieving the theoretical optimal [lan_communication-efficient_2020]. In this work, due to the different time scales, we generalize the CS procedure for problems involving a mixture of Bregman divergences. In addition, we point out that for problems with smooth objectives, Local SGD – which applies gradient steps locally but communicates only once in a while – has been studied under various settings [Zinkevich2010, stich2019local, Woodworth2020].
Primal-Dual Hybrid Gradient and its block variant. DCS [lan_communication-efficient_2020] is inspired by the Primal-Dual Hybrid Gradient (PDHG) algorithm [chambolle_first-order_2011, chambolle_ergodic_2016]. In this work, motivated by real-life settings where the costs – e.g. time and/or resources – needed for communication between different agents are different [Scaman2019], we further decompose the dual variables into blocks, and propose updating them at different frequencies. Thus, our algorithms can be viewed as multi-timescale variants of the PDHG. As a comparison, existing block-coordinate descent type of algorithms for the saddle point problem of interest, such as the Stochastic Primal-Dual Coordinate (SPDC) [pmlr-v37-zhanga15] and the Stochastic-PDHG (S-PDHG)[Alacaoglu2022, Chambolle2018Stochastic], update a random subset of blocks in each iteration $k$, where all blocks have strictly positive probability of being selected. The $O(1/k)$ rate of convergence, due to the randomness, is only shown for the expected objective value suboptimality (for SPDC) or expected duality gap (for S-PDHG). Nevertheless, in real applications, stochastic algorithms could potentially be less efficient, due to reasons such as random memory access [Sun2017] and potential overhead in computing sampling distributions [Csiba2018]. Although deterministic block coordinate descent for convex optimization has been shown to converge, such as under the cyclic updating rule [Sun2017, wright_coordinate_2015], to the best of our knowledge, the multi-timescale updating rule we propose is the first deterministic block updating rule for PDHG with separable duals, such that different blocks could be updated different numbers of times, and the duality gap converges deterministically at the optimal rate.
@@
Lower bounds on communication. In [ArjevaniShamir2015Communication], it is shown that for distributed convex optimization, $O(1/\epsilon)$ rounds of communication are needed for $1$-Lipschitz objectives, and $O(1/\sqrt{\mu\epsilon})$ rounds are needed when the objectives are also $\mu$-strongly convex. These lower bounds are achieved by splitting a “chain like” objective into two, each given to one agent. [Scaman2019] extends these results to a decentralized, network setting and shows the dependence of the lower bounds on the network diameter and communication delay. [Tyurin2024] provides lower bound and (nearly) optimal algorithm for a different setup, where distributed agents have stochastic first order oracles to the same smooth nonconvex objective, but computation and communication speeds are bounded and different for different edges and agents. Apart from the round complexity, [Santosh2020] shows a dimension-dependent lower bound on the bit-complexity of communication.
@@
In addition, motivated by distributed training in machine learning, communication lower and upper bounds have been established using function similarities [ArjevaniShamir2015Communication, pmlr-v32-shamir14, Zhu2016, pmlr-v151-tian22b, Kovalev2022, pmlr-v119-karimireddy20a, pmlr-v119-koloskova20a]: for instance, in (distributed) empirical risk minimization, the local loss functions have the same functional form but use different subsets of data, thereby inheriting the similarity in data. In [ArjevaniShamir2015Communication], function similarities are measured using norms of the differences in (sub)gradients (and Hessians if exist), and a communication round lower bound linear in this measure is shown for convex Lipschitz objectives and strongly convex objectives. Known algorithms that take advantage of function similarities usually require additional assumptions such as strong convexity and smoothness [Zhu2016, pmlr-v32-shamir14, pmlr-v37-zhangb15, pmlr-v151-tian22b, Kovalev2022, pmlr-v119-karimireddy20a, pmlr-v119-koloskova20a]. As pointed out in [ArjevaniShamir2015Communication], there is no known algorithm which achieve these communication round lower bounds for non-smooth convex objectives. In this work, we formalize the notion of function similarity for non-smooth convex objectives (Definition 2.1), and show that the communication round complexity for our (A)MT-GS indeed achieve these lower bounds, thereby answering [ArjevaniShamir2015Communication]’s open question positively.

### v2

# Multi-Timescale Primal Dual Hybrid Gradient with Application to Distributed Optimization

@@ Abstract
We propose two variants of the Primal Dual Hybrid Gradient (PDHG) algorithm for saddle point problems with block decomposable duals, hereafter called Multi-Timescale PDHG (MT-PDHG) and its accelerated variant (AMT-PDHG). Through novel mixtures of Bregman divergence and multi-timescale extrapolations, our MT-PDHG and AMT-PDHG converge under arbitrary updating rates for different dual blocks while remaining fully deterministic and robust to extreme delays in dual updates.
We further apply our (A)MT-PDHG, augmented with the gradient sliding techniques introduced in [27, 28], to distributed optimization. The flexibility in choosing different updating rates for different blocks allows a more refined control over the communication rounds between different pairs of agents, thereby improving the efficiencies in settings with heterogeneity in local objectives and communication costs. Moreover, with careful choices of penalty levels, our algorithms show linear and thus optimal dependency on function similarities, a measure of how similar the gradients of local objectives are. This provides a positive answer to the open question whether such dependency is achievable for non-smooth objectives [4].
@@ 1 Introduction
We study the following saddle point problem with block decomposable dual variables and objectives:
$\min_{X\in\mathcal{X}}\max_{y_{s}\in\mathbb{R}^{n_{s}},~s\in[S]}F(X)+\sum_{s=1}^{S}\left(\langle K_{s}X,y_{s}\rangle-R_{s}^{*}(y_{s})\right),$ ($\mathcal{P}_{b}$)
@@
where $\mathcal{X}\subset\mathbb{R}^{d}$ is a non-empty convex set and $F:\mathcal{X}\to\mathbb{R}$ is a convex function such that there exists $\mu,M\geq 0$,
$\frac{\mu}{2}\|X-X^{\prime}\|^{2}\leq F(X)-F(X^{\prime})-\langle F^{\prime}(X^{\prime}),X-X^{\prime})\rangle\leq M\|X-X^{\prime}\|,\quad\forall X,X^{\prime}\in\mathcal{X},$ (1)
@@
where $F^{\prime}:\mathcal{X}\to\mathbb{R}^{d}$ is a subgradient oracle, i.e. for each $X^{\prime}\in\mathcal{X}$, $F^{\prime}(X^{\prime})\in\partial F(X^{\prime})$ is a subgradient. For instance, when $\|F^{\prime}\|_{*}\leq M^{\prime}$, then $M=2M^{\prime}$ holds. [^1] This is because $F(X^{\prime})\geq F(X)+\langle F^{\prime}(X),X^{\prime}-X\rangle$, and so using Cauchy–Schwarz inequality $F(X)-F(X^{\prime})-\langle F^{\prime}(X^{\prime}),X-X^{\prime}\rangle\leq\langle F^{\prime}(X)-F^{\prime}(X^{\prime}),X-X^{\prime}\rangle\leq 2M\|X-X^{\prime}\|.$ For each $s\in[S]$, $K_{s}\in\mathbb{R}^{n_{s}\times d}$ is a matrix, $R_{s}:\mathbb{R}^{n_{s}}\to\overline{\mathbb{R}}$ is a proper, convex, and lower-semicontinuous function, and $R_{s}^{*}:\mathbb{R}^{n_{s}}\to\overline{\mathbb{R}}$ is its Fenchel conjugate, defined as $R_{s}^{*}(y_{s})=\sup_{y_{s}^{\prime}\in\mathbb{R}^{n_{s}}}\langle y_{s},y_{s}^{\prime}\rangle-R_{s}(y^{\prime}_{s})$.
Saddle point problems of the form ($\mathcal{P}_{b}$) with one block ($S=1$) have been widely studied due to its applications in various problems including linear programming [3], distributed optimization [27], and inverse problems such as image denoising [15, 14], to name a few. For $S>1$, stochastic algorithms which update random subsets of dual blocks have been proposed [13, 42].
In this work, building upon the Primal Dual Hybrid Gradient (PDHG) algorithms[15, 14], we propose Multi-Timescale Primal Dual Hybrid Gradient (MT-PDHG) and its accelerated variant (AMT-PDHG) where dual blocks are updated periodically at potentially different rates. Our algorithms achieve the following three properties.
  1. 1.
    Flexible. (A)MT-PDHG converge under arbitrary updating rates for the dual blocks.
  2. 2.
    Deterministic. The updating schedules for the dual blocks are deterministic and periodic.
  3. 3.
    Robust. The convergence rates depend on the average, instead of the maximum, of the updating rates, thereby are robust to extreme delays (i.e. large updating rates).
The motivation behind this work is distributed optimization, a branch of optimization where multiple agents, each having access to only partial information about the (global) objective, work together to solve the global problem. As an example, in distributed empirical risk minimization for machine learning, the global objective function is the sum of local loss functions, each depending on the local dataset which is only available to one agent [5, 11, 2, 4, 25]. Examples of other applications include power system control [32, 33], multi-robot system control [12, 33, 21, 44], and signal processing [9, 29, 40]. More precisely, we study the following distributed optimization problem
$\min_{x\in\overline{\mathcal{X}}}\sum_{v\in V}f_{v}(x)$ ($\mathcal{P}_{d}$)
where $V=[m]$ represents $m$ agents, $\overline{\mathcal{X}}\subset\mathbb{R}^{\overline{d}}$ is a nonempty, closed convex set, and $x^{*}\in\overline{\mathcal{X}}$ is an optimal solution to ($\mathcal{P}_{d}$).
From ($\mathcal{P}_{b}$) to lifted space reformulation of ($\mathcal{P}_{d}$). We consider a lifted space reformulation of ($\mathcal{P}_{d}$) – ($\mathcal{P}_{d}^{lift}$) in Section 4.1 – where agent $v$ maintains and updates $x_{v}$, a local version of the decision variable $x$. The objective then becomes $\sum_{v\in V}f_{v}(x_{v})$, with additional consensus constraints encouraging $x_{v}\approx x_{v^{\prime}}$ for all $v,v^{\prime}\in V$. The classical PDHG algorithm, when implemented in the distributed setting, requires communication only when dual variables are updated.
From ($\mathcal{P}_{d}$) and ($\mathcal{P}_{d}^{lift}$) to algorithm design goals for ($\mathcal{P}_{b}$). We target at distributed optimization under: (1) heterogeneity in local objectives and communication costs, and (2) communication bottlenecks with periodic connectivity constraints.
As an example of heterogeneity, in distributed empirical risk minimization, local loss functions are different due to randomness in the local datasets and variation of data distributions [2, 26, 43, 56, 48, 4]; the costs of communication could depend on factors such as the distance between agents, methods of communication, and amounts of data sent [8, 9, 41, 50, 51]. The heterogeneity makes it desirable to design algorithms allowing more flexible control over the numbers of communication rounds among different subsets of agents, which translates to the numbers of updates applied to each dual blocks. Existing block coordinate descent algorithms allow some level of control on this by using different selection probabilities for different blocks. However, stochastic algorithms could be impractical or inefficient due to factors such as unpredictability, random memory access [47], sampling overhead [18], and physical constraints on connectivity between agents [35].
We consider a “rate based design”, where communication is scheduled at different rates between different pairs of agents, deterministically. In practice, the rates could be either picked by users and/or subject to physical, bandwidth, and/or energy constraints [55]. This brings in one challenge: how to ensure the convergence of the algorithms under periodic communication at different rates. Another challenge, and also our third desideratum, is to ensure that the performance of the algorithms should not be influenced by stragglers (dual blocks which are updated infrequently), meaning that the algorithms should be robust to extreme values in the updating rates.
@@
Results. Our (A)MT-PDHG achieves the above three goals. To ensure convergence, we use careful mixtures of Bregman divergence and novel multi-timescale extrapolation. For the distributed optimization problem considered, to find an $\epsilon$-suboptimal solution, the complexities of our algorithms achieve optimal dependency on $\epsilon$: MT-PDHG needs $O(\overline{r}A/\epsilon)$ communication rounds and $O(\overline{r}/\epsilon^{2})$ subgradient steps for Lipchitz objectives, and AMT-PDHG needs $O(\overline{r}A/\sqrt{\epsilon\mu})$ communication rounds and $O(\overline{r}/(\epsilon\mu))$ subgradient steps if the objectives are also $\mu$-strongly convex. Here, $\overline{r}$ measures the “average rate of updates” for dual blocks, and $A$ measures similarities between (subgradients of) local functions. In fact, the linear dependency of communication rounds on $A$ is optimal [4], thereby providing a positive answer to the open question whether such dependency is achievable for non-smooth objectives [4].
@@
Numerical experiments for linear programming and support vector machine problem with regularized hinge losses confirm the effectiveness of our algorithms and demonstrate the above dependence on $\overline{r}$ and $A$.
@@ 1.1 Related works
Primal-Dual Hybrid Gradient and its block variant. Our algorithms are built upon the Primal-Dual Hybrid Gradient (PDHG) algorithms [14, 15]. For problems with block-decomposable duals, block-coordinate descent type of variants have been proposed, such as the Stochastic Primal-Dual Coordinate (SPDC) [57] and the Stochastic-PDHG (S-PDHG)[1, 13], to name a few. These algorithms update random subsets of the dual blocks at each iteration, where all blocks have strictly positive probability of being selected. Although the $O(1/k)$ rate of convergence still holds, due to the randomness, the convergence is only shown for the expected objective value suboptimality (for SPDC) or expected duality gap (for S-PDHG), and could have unsatisfactory performance due to large variance. Moreover, updating random subsets could potentially be inefficient – due to reasons such as random memory access [47] and potential overhead in computing sampling distributions [18] – or even impossible due to physical constraints – such as in distributed settings, blocks available for updates are subject to network connectivity and communication constraints [35].
Although deterministic block coordinate descent for convex optimization has been shown to converge, such as under the cyclic updating rule [47, 54], to the best of our knowledge, the multi-timescale updating rule we propose is the first deterministic block updating rule for PDHG with separable duals, such that different blocks could be updated different numbers of times, and the duality gap converges deterministically at the optimal rate.
As the dual blocks are not updated at each global iteration, they introduce ”outdated” information to the dynamics. Most existing asynchronous optimization algorithms achieve convergence rates which depend on the maximum delay [37, 30, 38], with the exception of [17, 6], whose convergence rates depend on the average and the quantiles of delays, respectively. However, [17, 6] are designed for convex optimization. As a comparison, our (A)MT-PDHG are designed for saddle point problems and have convergence rates which depend on the averages instead of maximum of the updating rates, making them robust against extreme delays in the dual updates.
@@
Non-smooth distributed optimization. Since the seminal works [9, 49], numerous algorithms have been proposed for non-smooth distributed optimization under various settings, and we refer readers to surveys such as [5, 34, 24]. For the function class of Lipschitz, non-smooth, convex objectives, most of these algorithms fall into the following two categories: subgradient based and dual based [27]. Subgradient based algorithms such as the incremental gradient method [10], decentralized subgradient method [36], and the dual averaging [19] usually require $O(1/\epsilon^{2})$ rounds of communication, each followed by one gradient step. Within the function class, this achieves the optimal subgradient oracle complexity, but is suboptimal with respect to the communication rounds: as proven by [4, 41], the communication rounds needed is $O(1/\epsilon)$.
@@
As a comparison, dual based algorithms, which dualize the consensus constraints, usually have better communication complexity: $O(1/\epsilon)$ rounds are needed for distributed ADMM [7, 52] and the decentralized communication sliding (DCS) [27], as examples. However, each round of communication is followed by optimization of Lagrangians or proximal updates, performed locally by each agent. To make the overall algorithm first-order, [27] proposes the Communication Sliding (CS) procedure, which approximates the proximal updates through $O(1/\epsilon)$ steps of (local) mirror descent, thereby achieving the $O(1/\epsilon^{2})$ subgradient oracle complexity. The CS procedure has roots in the gradient sliding technique [28], which can save gradient computation for the smooth component when the objective involves a smooth and a non-smooth component.
@@
For the class of strongly convex objectives, DCS can be accelerated, needing $O(1/\sqrt{\epsilon})$ rounds of communication and $O(1/\epsilon)$ gradient steps in total, both achieving the theoretical optimal [27]. In this work, due to the different time scales, we generalize the CS procedure for problems involving a mixture of Bregman divergences. In addition, we point out that for problems with smooth objectives, Local SGD – which applies gradient steps locally but communicates only once in a while – has been studied under various settings [58, 46, 53].
@@
Lower bounds on communication. In [4], it is shown that for distributed convex optimization, $O(1/\epsilon)$ rounds of communication are needed for $1$-Lipschitz objectives, and $O(1/\sqrt{\mu\epsilon})$ rounds are needed when the objectives are also $\mu$-strongly convex. These lower bounds are achieved by splitting a “chain like” objective into two, each given to one agent. [41] extends these results to a decentralized, network setting and shows the dependence of the lower bounds on the network diameter and communication delay. [50] provides lower bound and (nearly) optimal algorithm for a different setup, where distributed agents have stochastic first order oracles to the same smooth nonconvex objective, but computation and communication speeds are bounded and different for different edges and agents. Apart from the round complexity, [51] shows a dimension-dependent lower bound on the bit-complexity of communication.
@@
In addition, motivated by distributed training in machine learning, communication lower and upper bounds have been established using function similarities [4, 43, 2, 48, 26, 22, 23]: for instance, in (distributed) empirical risk minimization, the local loss functions have the same functional form but use different subsets of data, thereby inheriting the similarity in data. In [4], function similarities are measured using norms of the differences in (sub)gradients (and Hessians if exist), and a communication round lower bound linear in this measure is shown for convex Lipschitz objectives and strongly convex objectives. Known algorithms that take advantage of function similarities usually require additional assumptions such as strong convexity and smoothness [2, 43, 56, 48, 26, 22, 23]. As pointed out in [4], there is no known algorithm which achieve these communication round lower bounds for non-smooth convex objectives. In this work, we formalize the notion of function similarity for non-smooth convex objectives (Definition 4.1), and show that the communication round complexity for our (A)MT-PDHG indeed achieve these lower bounds, thereby answering [4]’s open question positively.